Please fill the table for your shift with this information:

  • PROGRAM DONE: please write the type of measurements you have done, with reference to run number range and to logbook pages
  • BEAM INFO: note periods with no beam, report related problems
  • ASIDE ACTIVITIES: please add information on aside activities such as debugging… put references to logbook pages
  • OBSERVED PROBLEMS: add Meaningful information of problems encountered in your shift

Shift Date 16.7.2012
Shift Time 16-24
Shifter Names Daniele, John
Program Done Found center of PbF2 crystal; started runs to compare polarizer orientations. Conclusion: saturation of the narrow Cerenkov pulses distorts the ADC means. Lowered HV from 700V to 600V.
Beam Info -180 GeV pion
Aside Activities  
Observed Problems  

Shift Date 17.7.2012
Shift Time 0-8
Shifter Names Antonio, Gabriella
Program Done Trying to understand the polirizer filter in the PBF2 xtal matrix. Trying to set electron beam, but the rate was very low. Logbook pages 20-23 ; Runs from 6700 to 6706
Beam Info H8C.012 +80 GeV electron
Aside Activities none
Observed Problems mismatch of the T1&T2 numeber as provided out of the trigger box and that made with NIM coincidence using discriminator signals

Shift Date 17.7.2012
Shift Time 8-16
Shifter Names Michele C., Evelin
Program Done HV scan. Removed PS detector to start shower depth study, HV scan restarted again due to the large effect of PS detector removal. Ilias has tried to fix the electron beam. Trying to reproduce the polarization effect observed in 2010.
Beam Info H8C.012 +80 GeV electron
Aside Activities none
Observed Problems One hour without beam

Shift Date 17.7.2012
Shift Time 16-24
Shifter Names Alan Sill, Richard Wigmans
Program Done Found reason for lack of difference between orientation directions for polarizers for PbF2 crystal: it turns out that this crystal scrambles polarization when viewed along its long axis. Switched to negative polarity for beams and established good pion and electron beams. Also verified XY distributions for both veto and non-veto events. (Note "veto" means its inverse here, i.e., events for particles that pass throughthe hole in the veto counter.) Changed angle of crystal to 30 degrees and started program to repeat depth dependence of BSO signal.
Beam Info pi- 180 GeV, e- 80 GeV
Aside Activities Studied polarization properties of PbF2 crystal
Observed Problems none

Shift Date 18.7.2012
Shift Time 0-8
Shifter Names Antonio, Gabriella
Program Done BSO xtal polarization measurements. Scan as function of the depth of the material (lead) positioned upstream the crystal. No veto and no preshower detectors in the daq. Favorable orientation of the polarizer filter: scan from 0 X_0 to 10 X_0 (Runs from 6759 to 6770; logbook pages 36-38). Scan with polarizer in unfavorable orientation,scan from 4 Xo to 10 Xo (runs from 6771 to 6779, logbook page 39)
Beam Info good beam all the night except 20 min 5 am; MD started at 7 am; run 6779 only 5k evt
Aside Activities none
Observed Problems none

Shift Date 18.7.2012
Shift Time 16-24
Shifter Names guido,nural
Program Done restart after md at 19:00; problems with the beam dumps; e- 80 gev beam file lost. New e- 80 gev file : H729. Re-do all X0 scan due to the new beam setup. Data taking restarted at 11 pm. 4/3/2 X0 unfovorable
Beam Info New e- beam file H729
Aside Activities problem with pedestal of chan S3L solved by Agostino who replaced the amplificator chip
Observed Problems beam dumps blocked; Ilias called

Shift Date 19.7.2012
Shift Time 0-8
Shifter Names Silvia, Evelin
Program Done BSO polarisation scan in X0, unfavorable polarizer (pag 47 logbook). Redone measurements of yesterday night because beam file is changed
Beam Info New e- beam file H729
Aside Activities  
Observed Problems  

Shift Date 19.7.2012
Shift Time 8-16
Shifter Names Antonio/Nural
Program Done BSO xtal polarization measurements. Completed the X0 scan with polarizer in unfavorable orientation (pag 47 of the logbook). X0 scan with polarizer in favorable orientation from run 6817 to run 6830 (pag 48 of the logbook).
Beam Info 80GeV e- beam file H729, no beam from 7:30 to 09:30; from 11:00 to 11:40 and from 12:10 to 13:30.
Aside Activities none
Observed Problems none

Shift Date 19.7.2012
Shift Time 16-24
Shifter Names John, Daniele
Program Done Richard setted beamfiles for different electron energies, moved the table with Pavia 2011 newdream in the beam, started to look at signals
Beam Info  
Aside Activities  
Observed Problems  

Shift Date 20.7.2012
Shift Time 0-8
Shifter Names Silvia, Evelin
Program Done Found edges of newdream with pion beam, setted HV for each tower with e beam 80 Gev in order to have ~ 800 ADC cnts above pedestal. Seen that the centre of each tower found with pion beam is different (in particular on X coordinate wrt to teh electron beam (dept of shower and angle of calorimeter..) we did again a scan in position. Found new coordinates for towers and setted HV properly. Started calibration with 20k events in each tower
Beam Info pi beam + 80GeV e- beam file H729
Aside Activities  
Observed Problems  

Shift Date 20.7.2012
Shift Time 8-16
Shifter Names Antonio & Nural
Program Done Started the energy scan in T3. We have discovered a problem with response: we observe non linearity in the Cerenkov at high energy, due to the C3 phototube saturation. Moving the C3 HV from 830V to 700V and start a new energy scan (pag. 56-58 of the logbook). The new energy scan has been stopped because we discovered that there is a linearity problem (counts/GeV) in the low energy region (see again pag 56-58 of the logbook): the problem seems to be related to the ADC (see pag 56-59 of the logbook).
Beam Info 80GeV e- beam file H729 for calibration; see the twiki page for the beam files of the energy scan.
Aside Activities none
Observed Problems see the "program done"

Shift Date 20.7.2012
Shift Time 16-24
Shifter Names Tomasso, Alan
Program Done Energy scan , no mixers. Results show detector is very linear in Cerenkov signal also. Still using scope since ADC readout problem is not understood. Proceeded to test C3, S3 with mixers also.
Beam Info e- various energies
Aside Activities  
Observed Problems DAQ frequently stops due to mismatch between number of scope events and number of VME events when running at low rates

Shift Date 21.7.2012
Shift Time 0-8
Shifter Names Sehwook, Gabri
Program Done Energy scan with electron beam on tapered base, tapered with independent power supply for last dynode and standard base. Run 6982 to 7029, logbook pages 69-74
Beam Info beam stable all time
Aside Activities none
Observed Problems none

Shift Date 21.7.2012
Shift Time 8-16
Shifter Names Guido, Silvietta
Program Done Test ADC after Alessandro fix; apparently the linearity has been recovered but also very funny behaviour on some channel (saturation....). Back to Oscilloscope. Start testing the light mixers 4/2 cm.
Beam Info -80 Gev electrons for the mixer tests; good beam during the whole shift
Aside Activities light mixer installation
Observed Problems stop of DAQ during the run; see Alan comments

Shift Date 21.7.2012
Shift Time 16-24
Shifter Names Alan, Tommaso
Program Done continue light mixer tests
Beam Info stable beam
Aside Activities  
Observed Problems  

Shift Date 22.7.2012
Shift Time 0-8
Shifter Names Sehwook, Gabri
Program Done Completed light mixer test; Uninstalled light mixer and back to normal configuration; rotated table at 0 degrees, and found module position (logbook page 82); Started data taking for electron to pion separation based on the different time structure of the signal. 80 GeV electron and pion data taken, 30 GeV pion ongoing (runs 7094 - 7097; logbook page 83)
Beam Info stable beam
Aside Activities Measured lead plates (see logbook page 84)
Observed Problems none

Shift Date 22.7.2012
Shift Time 8-16
Shifter Names Laura, Silvietta
Program Done ended data taking for e-pi separation run 7098 e- 30 GeV - run 7099 e/pai 125 GeV (pg. 80).Polarization studies (pg 85)
Beam Info stable beam
Aside Activities  
Observed Problems none

Shift Date 22.7.2012
Shift Time 16-24
Shifter Names Alan, Franco
Program Done  
Beam Info  
Aside Activities tapered bases mounted on each tower of newdream
Observed Problems oscilloscope not working with 4 channels

Shift Date 23.7.2012
Shift Time 0-8
Shifter Names Silvia- Tommaso
Program Done equalised gains of the 4 towers wrt to T3 that we tested on this week (now tapered bases for each PMT). Equalization done after the stage of linear fan-in used to sum signals for the four towers (we have only two scope signals available). Equalisation done also after the sum of the four signals. Energy scan started with beam in the centre of newdream
Beam Info energy scan, elecrons
Aside Activities  
Observed Problems  

Shift Date 23.7.2012
Shift Time 8-16
Shifter Names Richard, Daniele
Program Done module scan in 2 cm steps
Beam Info 80 gev e-
Aside Activities start Pisa modules install
Observed Problems  

Shift Date 23.7.2012
Shift Time 16-24
Shifter Names Guido, Franco
Program Done Start Pisa Module analysis. Find module position; find position of towers. Fast energy scan on C3 S3
Beam Info 80 gev e-
Aside Activities end of Pisa module installation
Observed Problems beam dump blocked after the access

Shift Date 24.7.2012
Shift Time 0-8
Shifter Names Laura,Tommaso
Program Done equalization Pisa Module (logbook pg.98) - energy scan Pisa module center (e-) : 80 - 100 - 125 -150 GeV
Beam Info no/instable beam from 5.15; from 5.54 no beam
Aside Activities  
Observed Problems DQA problem: beam was not detected. Bob and Gabriella solved it. About 2 h were lost in data acquisition

Shift Date 24.7.2012
Shift Time 8-16
Shifter Names John, Evelin
Program Done Continued the energy scan Pisa module center (e-) : 50 GeV, 30 GeV, 20 GeV - Started the position scan of Pisa module with 80 GeV electrons (Logbook p.98-100)
Beam Info no beam before 8:50;
Aside Activities  
Observed Problems  

Shift Date 24.7.2012
Shift Time 16-24
Shifter Names Daniele,Franco L.
Program Done Second copper module from Pisa. Position scan. Energy scan for tower 3. Logbook p. 101-104
Beam Info  
Aside Activities  
Observed Problems  

Shift Date 24.7.2012
Shift Time 0-8
Shifter Names Silvia Tommaso
Program Done equalization towers second Cu module, Energy scan, beam in the centre
Beam Info energy scan
Aside Activities ntuple analysed offline, no problems encountered
Observed Problems some runs (7368) with empty events in the scope. Some jittering coming out from the adder on S channel

Shift Date 25.7.2012
Shift Time 16-24
Shifter Names Laura Franco
Program Done -
Beam Info no beam - other problems after machine development
Aside Activities Bob cardini gabriella worked to solve ADC problems.(pg. 107-109 logbook)
Observed Problems we'll continue to use oscilloscope!

Shift Date 26.07.12
Shift Time 00:00-08:00
Shifter Names Daniele, Fabrizio
Program Done Nothing ! We weren't able to setup a good 60 GeV pion beam file to derive the low energy electron beam files
Beam Info no beam till 02:00
Aside Activities modified the macro to view the monitor histograms; included ch.3 and ch.4 of the digital scope
Observed Problems no comment

Shift Date 26.07.12
Shift Time 8-16
Shifter Names guido, tommaso
Program Done  
Beam Info no beam 10 till 16
Aside Activities  
Observed Problems beam off

Shift Date 26.07.12
Shift Time 16-24
Shifter Names Silvia, Alessandro
Program Done Cu_2 module, T3 energy scan at low energy beam
Beam Info low energy files, wobbling changed to 60 GeV
Aside Activities Ilias changed some beam parameter and doubled beam rate
Observed Problems  

Shift Date 26.07.12
Shift Time  
Shifter Names  
Program Done  
Beam Info  
Aside Activities  
Observed Problems  

Shift Date 26.07.12
Shift Time 0-8
Shifter Names Sehwook Laura
Program Done energy scan in center module (logbook pg. 112-113)
Beam Info stable beam
Aside Activities none
Observed Problems the indicated scope setting at 8 GeV are probably not correct. See the logbook

Shift Date 27.07.12
Shift Time 8-16
Shifter Names John, Evelin
Program Done Completed energy scan in center module (logbook p.112), ene: 4 GeV- 8 GeV. Done scan position (with ele 20 GeV) (logbook p. 115). Started energy scan in T3, ene : 20 GeV, 30 GeV, 40 GeV, 50 GeV (logbook p. 114)
Beam Info  
Aside Activities  
Observed Problems  

Shift Date 27.07.12
Shift Time 16-24
Shifter Names  
Program Done  
Beam Info  
Aside Activities  
Observed Problems  

Shift Date 28.07.12
Shift Time 0-8
Shifter Names Alan Laura
Program Done Pb module: 30 geV at center of each tower - energy scan tow3 - energy scan module center (logbook pg 118-121
Beam Info stable beam
Aside Activities none
Observed Problems  

Shift Date 26.07.12
Shift Time  
Shifter Names  
Program Done  
Beam Info  
Aside Activities  
Observed Problems  

Shift Date 28.07.12
Shift Time 16-24
Shifter Names Fabrizio Laura
Program Done Cu-Al module Energy scan tow3 - calibration of leakage counters started- the calorimeter slope has been removed (logbook pg. 131-133)
Beam Info 1 hour missing
Aside Activities leakages counters installed for calibration
Observed Problems  

Shift Date 26.07.12
Shift Time  
Shifter Names  
Program Done  
Beam Info  
Aside Activities  
Observed Problems  

Shift Date 26.07.12
Shift Time  
Shifter Names  
Program Done  
Beam Info  
Aside Activities  
Observed Problems  

Shift Date 26.07.12
Shift Time  
Shifter Names  
Program Done  
Beam Info  
Aside Activities  
Observed Problems  

-- GabriellaGaudio - 16-Jul-2012

Edit | Attach | Watch | Print version | History: r46 < r45 < r44 < r43 < r42 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r46 - 2012-07-29 - LauraLaRotonda
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    DREAM All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback